# Joint Diffusion Models in Continual Learning

### v2

@@ Abstract
In this work, we introduce JDCL – a new method for continual learning with generative rehearsal based on joint diffusion models. Neural networks suffer from catastrophic forgetting defined as abrupt loss in the model’s performance when retrained with additional data coming from a different distribution. Generative-replay-based continual learning methods try to mitigate this issue by retraining a model with a combination of new and rehearsal data sampled from a generative model. In this work, we propose to extend this idea by combining a continually trained classifier with a diffusion-based generative model into a single – jointly optimized neural network. We show that such shared parametrization, combined with the knowledge distillation technique allows for stable adaptation to new tasks without catastrophic forgetting. We evaluate our approach on several benchmarks, where it outperforms recent state-of-the-art generative replay techniques. Additionally, we extend our method to the semi-supervised continual learning setup, where it outperforms competing buffer-based replay techniques, and evaluate, in a self-supervised manner, the quality of trained representations.
@@ 1 Introduction
Figure 1: Further training of a classifier using data generated from the diffusion model trained with the same dataset harms the classifier performance (blue line). However, joint modeling, especially in combination with knowledge distillation significantly limits this degradation and retains almost the initial performance. We report accuracy averaged across three different seeds.
To visualize the main benefit of this approach, we conduct a simple experiment that highlights the drawbacks of recent techniques. We train the classifier and our joint-diffusion model (ensuring that both models are of similar sizes) on the CIFAR10 dataset until convergence (for 100,000 optimizer steps) [^1] Details on this experiment are provided in the Appendix. Then, we continue to optimize the models, but instead of using real data, we sample data from the diffusion model. As presented in Fig. 1, the performance of the classifier (blue line), drastically decreases as the effect of the rehearsal with generated samples. At the same time, the performance of our joint diffusion (yellow) converges after a much smaller drop in performance. Furthermore, the addition of knowledge distillation significantly decreases the degradation of both models, while keeping the trend of the joint diffusion model outperforming the classifier in terms of final performance and stability. We conclude that a combination of joint modeling and knowledge distillation is a key for knowledge retention in generative replay-base CL.
@@ 2 Related Work
The increasing size of machine learning models, and hence the increased computational cost of their training highlights the need for effective methods for model updates. Therefore, we can observe a growing number of continual learning techniques [45, 43] that aim to mitigate the problem of catastrophic forgetting [16], without reducing the ability to adjust to new data. CL methods can be categorized into three groups: (1) Regularization-based approaches such as EWC [27], LwF [35] reduce the interference from the new tasks on the parts of the neural network that are crucial for the previous ones. (2) Architecture-based methods (e.g.RCL [75], PNN [53]) directly build different model versions for different tasks. Finally (3) Replay techniques (e.g. [48, 1]) use a limited number of past samples as rehearsal examples, by mixing them together with new data samples.
@@
A significant drawback of replay approaches is that storing past samples requires a growing memory buffer. Therefore, Shin et al. [58], proposed to replace it with a Generative Adversarial Network [20]. This idea was further extended to VAEs [66, 42], GMMs [51], or Diffusion-based generative models [18, 8]. In [33] authors provide a summary of different generative replay techniques. Apart from different architectures, similar to our idea, an interesting approach is proposed in Replay Through Feedback (RTF) [66], where authors combine the variational autoencoder with a classifier. This improves the performance of the final solution, which is in line with results presented in [64] and [40], where authors show that additional reconstruction or generative loss added to the classifier alleviates catastrophic forgetting.
@@
Several methods explore the potential of using the latent representations within the U-Net model used as a denoising decoder in DDGM. In particular, Kwon et al. [30], show that the bottleneck of the U-Net dubbed H-space, can be used as a semantic latent space for image manipulation. This idea is further extended in [46]. Similarly, some techniques employ the H-space in downstream tasks outside of the generative modeling. Authors of [2], [65] and [50], show that such features can be used for image segmentation, in [38] Luo et al. extend this idea to image correspondence, while  [11] use H-features were used for classification. In this work, we relate to the utilization of H-space by employing it in a continually trained joint diffusion model with a classifier.
@@
Semi-supervised methods confront the problem of training models on large datasets where only a few samples have associated annotations [61, 76]. Recently, most of the best-performing semi-supervised methods belong to the family of hybrid methods that combine consistency regularization and pseudo-labeling  [4, 29, 61, 76, 78]. The main idea behind consistency regularization is an assumption that the model should yield similar outputs when fed perturbed versions of the same image or when its weights are slightly perturbed [31, 54, 63, 41]. On the other hand, pseudo-labeling techniques exploit the concept of using synthetic labels for unlabeled data. More precisely, this involves utilizing ”hard” labels (i.e., the arg max of the model’s output) for which the predicted class probability surpasses a predetermined threshold [74, 7, 71, 47, 32, 76]. Apart from consistency regularization and pseudo-labeling, a third group of SSL methods employs additional deep generative models  [76] such as Variational Auto-Encoder (VAE) [34, 44, 14, 39, 26], or Generative Adversarial Networks (GAN)  [55, 62, 73, 17, 37], to learn the underlying class-conditional distribution of partially labeled data.
Despite their proliferation, as noted by [25], the standard semi-supervised approaches do not address the problem of learning in a continual setup. Consequently, a new line of research called continual semi-supervised learning emerged with the goal of addressing the challenges of the combination of those two domains. To that end, [70] introduced ORDisCo, which utilizes a conditional GAN with a classifier to learn from partially labeled data in an incremental setup. On the other hand, [5] proposed CCIC, a method that enforces consistency by contrasting samples among different classes and tasks. In contrast, the recent state-of-the-art method called NNCSL [25] leverages the nearest-neighbor classifier to learn the current task, while also retaining the relevant information from the previous tasks. In this work, we adapt the mechanism of pseudo-labeling and consistency regularization, which we combine with our joint diffusion model to combat the problems of semi-supervised continual learning.
@@ 3.2 Joint Diffusion Model
We follow the parametrization of joint diffusion models introduced in [11], where features from the different levels of the UNet architecture are pooled into a single vector $z$ used for classification. Let $p_{\theta}$ be the denoising model, which is of the UNet architecture, with parameters $\theta$. As such, the denoising model consists of the encoder $e_{\nu}$ and the decoder $d_{\psi}$ with parameters $\nu$ and $\psi$ respectively, such that $\theta=\{\nu,\psi\}$. The encoder takes input $x_{t}$ and returns a set of tensors $\mathcal{Z}_{t}=e_{\nu}(x_{t})$ that is, a set of representation tensors $\mathcal{Z}_{t}=\{z_{t}^{1},z_{t}^{2}\dots z_{t}^{n}\}$ obtained from each depth level of the UNet architecture. The decoder reconstructs the denoised sample $x_{t-1}$, from the set of representations, i.e., $x_{t-1}=d_{\psi}(\mathcal{Z}_{t})$. Additionally to the standard diffusion denoiser, we propose to pool the representation set into a single vector $z_{t}=f(\mathcal{Z}_{t})$ via average pooling. Finally, we introduce classifier $g_{\omega}$, with parameters $\omega$ the final part of the model that predicts target class $\hat{y}=g_{\omega}(z_{t})$.
@@
ϵ^=bold-^bold-italic-$\boldsymbol{\hat{\epsilon}}=$ $d_{\psi}(\{z_{t}^{1},z_{t}^{2}\dots z_{t}^{n}\}).$ (5)
@@
which corresponds to the cross-entropy loss, where $y$ represents the target class, $\varphi$ is a vector of probabilities generated by the classifier $g_{\omega}(e_{\nu}(x_{0}))$ and $\mathds{1}[y=k]$ is the indicator function that is $1$ if $y$ equals $k$ and 00 otherwise. The final loss function that we optimize with a single optimizer over parameters $\{\nu,\psi,\omega\}$ is:
@@ 4 Method
Figure 2: Overview of JDCL. To adapt the global model to task $\tau+1$, we first train a local joint diffusion model on task $\tau+1$ only. We then use the global model and the local model to generate a synthetic dataset consisting of samples from tasks $11,2\dots\tau+$ 1. Finally, we fine-tune the global model with the generated dataset and a combination of joint diffusion and knowledge distillation losses.
@@
One of the crucial problems in continually trained neural networks is the plasticity-stability dilemma. As presented in Fig 1, we can easily achieve high stability of our continually trained model through a self-rehearsal mechanism combined with knowledge distillation. However, to improve the plasticity, we relate to the two-stage training [10, 19] where we first fine-tune the copy of the current model using only newly available data which results in perfect plasticity. Then, we distill the adapted model with the previous one into a new global model as presented in Fig. 2.
@@ 5.1 Experimental Setup
We evaluate our method on three commonly used datasets – CIFAR10, CIFAR100 [28] and ImageNet100 [12]. The first two of those datasets contain 50000 training and 10000 test images of size $32\times 32$ divided into 10 or 100 classes. ImageNet100 on the other hand contains 120000 train images in $64\times 64$ resolution from 100 different classes. For the class-incremental setup, we split the CIFAR10 and ImageNet100 datasets into 5 equal tasks and the CIFAR100 dataset into 5 or 10 equal tasks. For semi-supervised setups, we train our models on two different ratios of labeled data in the dataset i.e. 0.8% and 5%. Those ratios correspond to 4 and 25 labelled examples per class in CIFAR100, and to 40 and 250 samples for CIFAR10.

### v3

@@ Abstract
In this work, we introduce JDCL – a new method for continual learning with generative rehearsal based on joint diffusion models. Neural networks suffer from catastrophic forgetting defined as an abrupt loss in the model’s performance when retrained with additional data coming from a different distribution. Generative-replay-based continual learning methods try to mitigate this issue by retraining a model with a combination of new and rehearsal data sampled from a generative model. In this work, we propose to extend this idea by combining a continually trained classifier with a diffusion-based generative model into a single – jointly optimized neural network. We show that such shared parametrization, combined with the knowledge distillation technique, allows for stable adaptation to new tasks without catastrophic forgetting. We evaluate our approach on several benchmarks, where it outperforms recent state-of-the-art generative replay techniques. Additionally, we extend our method to the semi-supervised continual learning setup, where it outperforms competing buffer-based replay techniques, and evaluate, in a self-supervised manner, the quality of trained representations.
@@ 1 Introduction
Figure 1: Further training of a classifier using data generated from the diffusion model trained with the same dataset harms the classifier’s performance (blue line). However, joint modeling, especially in combination with knowledge distillation, significantly limits this degradation and retains almost the initial performance. We report accuracy averaged across three different seeds.
To visualize the main benefit of this approach, we conduct a simple experiment that highlights the drawbacks of recent techniques. We train the classifier and our joint-diffusion model (ensuring that both models are of similar sizes) on the CIFAR10 dataset until convergence (for 100,000 optimizer steps) [^1] Details on this experiment are provided in the Appendix. Then, we continue to optimize the models, but instead of using real data, we sample data from the diffusion model. As presented in Fig. 1, the performance of the classifier (blue line), drastically decreases as the effect of the rehearsal with generated samples. At the same time, the performance of our joint diffusion (yellow) converges after a much smaller drop in performance. Furthermore, the addition of knowledge distillation significantly decreases the degradation of both models, while keeping the trend of the joint diffusion model outperforming the classifier in terms of final performance and stability. We conclude that a combination of joint modeling and knowledge distillation is key for knowledge retention in generative replay-base CL.
@@ 2 Related Work
The increasing size of machine learning models, and hence the increased computational cost of their training highlights the need for effective methods for model updates. Therefore, we can observe a growing number of continual learning techniques [47, 45] that aim to mitigate the problem of catastrophic forgetting [17], without reducing the ability to adjust to new data. CL methods can be categorized into three groups: (1) Regularization-based approaches such as EWC [28], LwF [36] reduce the interference from the new tasks on the parts of the neural network that are crucial for the previous ones. (2) Architecture-based methods (e.g.RCL [78], PNN [55]) directly build different model versions for different tasks. Finally (3) Replay techniques (e.g. [50, 1]) use a limited number of past samples as rehearsal examples, by mixing them with new data samples.
@@
A significant drawback of replay approaches is that storing past samples requires a growing memory buffer. Therefore, Shin et al. [60] proposed to replace it with a Generative Adversarial Network [21]. This idea was further extended to VAEs [69, 44], GMMs [53], or Diffusion-based generative models [19, 9]. In [34] authors provide a summary of different generative replay techniques. Apart from different architectures, similar to our idea, an interesting approach is proposed in Replay Through Feedback (RTF) [69], where authors combine the variational autoencoder with a classifier. This improves the performance of the final solution, which is in line with results presented in [67, 42], where authors show that additional reconstruction or generative loss added to the classifier alleviates catastrophic forgetting.
@@
Several methods explore the potential of using the latent representations within the U-Net model used as a denoising decoder in DDGM. In particular, Kwon et al. [31] show that the bottleneck of the U-Net dubbed H-space can be used as a semantic latent space for image manipulation. This idea is further extended in [48]. Similarly, some techniques employ the H-space in downstream tasks outside of generative modeling. Authors of [2], [68] and [52], show that such features can be used for image segmentation, in [39] Luo et al. extend this idea to image correspondence, while  [12] use H-features for classification. In this work, we relate to the utilization of H-space by employing it in a continually trained joint diffusion model with a classifier.
@@
Semi-supervised methods confront the problem of training models on large datasets where only a few samples have associated annotations [63, 80]. Recently, most of the best-performing semi-supervised methods belong to the family of hybrid methods that combine consistency regularization and pseudo-labeling  [4, 30, 63, 80, 82]. The main idea behind consistency regularization is an assumption that the model should yield similar outputs when fed perturbed versions of the same image or when its weights are slightly perturbed [32, 56, 66, 43]. On the other hand, pseudo-labeling techniques exploit the concept of using synthetic labels for unlabeled data. More precisely, this involves utilizing "hard" labels (i.e., the arg max of the model’s output) for which the predicted class probability surpasses a predetermined threshold [77, 7, 74, 49, 33, 80]. Apart from consistency regularization and pseudo-labeling, a third group of SSL methods employs additional deep generative models  [80] such as Variational Auto-Encoder (VAE) [35, 46, 15, 40, 27], or Generative Adversarial Networks (GAN)  [57, 64, 76, 18, 38], to learn the underlying class-conditional distribution of partially labeled data.
Despite their proliferation, as noted by [26], the standard semi-supervised approaches do not address the problem of learning in a continual setup. Consequently, a new line of research called continual semi-supervised learning emerged to address the challenges of the combination of those two domains. To that end, [73] introduced ORDisCo, which utilizes a conditional GAN with a classifier to learn from partially labeled data in an incremental setup. On the other hand, [5] proposed CCIC, a method that enforces consistency by contrasting samples among different classes and tasks. In contrast, the recent state-of-the-art method called NNCSL [26] leverages the nearest-neighbor classifier to learn the current task, while also retaining the relevant information from the previous tasks. In this work, we adapt the mechanism of pseudo-labeling and consistency regularization, which we combine with our joint diffusion model to combat the problems of semi-supervised continual learning.
@@ 3.2 Joint Diffusion Model
We follow the parametrization of joint diffusion models introduced in [12], where features from the different levels of the UNet architecture are pooled into a single vector $z$ used for classification. Let $p_{\theta}$ be the denoising model, which is of the UNet architecture, with parameters $\theta$. As such, the denoising model consists of the encoder $e_{\nu}$ and the decoder $d_{\psi}$ with parameters $\nu$ and $\psi$ respectively, such that $\theta=\{\nu,\psi\}$. The encoder takes input $x_{t}$ and returns a set of tensors $\mathcal{Z}_{t}=e_{\nu}(x_{t})$ that is, a set of representation tensors $\mathcal{Z}_{t}=\{z_{t}^{1},z_{t}^{2}\dots z_{t}^{n}\}$ obtained from each depth level of the UNet architecture. The decoder reconstructs the denoised sample $x_{t-1}$, from the set of representations, i.e., $x_{t-1}=d_{\psi}(\mathcal{Z}_{t})$. Additionally to the standard diffusion denoiser, we propose to pool the representation set into a single vector $z_{t}=f(\mathcal{Z}_{t})$ via average pooling. Finally, we introduce classifier $g_{\omega}$, with parameters $\omega$ - the final part of the model that predicts target class $\hat{y}=g_{\omega}(z_{t})$.
@@
$\boldsymbol{\hat{\epsilon}}$= $d_{\psi}(\{z_{t}^{1},z_{t}^{2}\dots z_{t}^{n}\}).$ (5)
@@
which corresponds to the cross-entropy loss, where $y$ represents the target class, $\varphi$ is a vector of probabilities generated by the classifier $g_{\omega}(e_{\nu}(x_{0}))$ and $\mathds{1}[y=k]$ is the indicator function that is $1$ if $y$ equals $k$ and 0 otherwise. The final loss function that we optimize with a single optimizer over parameters $\{\nu,\psi,\omega\}$ is:
@@ 4 Method
Figure 2: Overview of JDCL. To adapt the global model to task $\tau+1$, we first train a local joint diffusion model on task $\tau+1$ only. We then use the global model and the local model to generate a synthetic dataset consisting of samples from tasks $1,2\dots\tau+1$. Finally, we fine-tune the global model with the generated dataset and a combination of joint diffusion and knowledge distillation losses.
@@
One of the crucial problems in continually trained neural networks is the plasticity-stability dilemma. As presented in Fig 1, we can easily achieve high stability of our continually trained model through a self-rehearsal mechanism combined with knowledge distillation. However, to improve the plasticity, we relate to the two-stage training [11, 20] where we first fine-tune the copy of the current model using only newly available data, which results in perfect plasticity. Then, we distill the adapted model with the previous one into a new global model as presented in Fig. 2.
@@ 5.1 Experimental Setup
We evaluate our method on three commonly used datasets – CIFAR10, CIFAR100 [29] and ImageNet100 [13]. The first two of those datasets contain 50000 training and 10000 test images of size $32\times 32$ divided into 10 or 100 classes. ImageNet100 on the other hand contains 120000 train images in $64\times 64$ resolution from 100 different classes. For the class-incremental setup, we split the CIFAR10 and ImageNet100 datasets into 5 equal tasks and the CIFAR100 dataset into 5 or 10 equal tasks. For semi-supervised setups, we train our models on two different ratios of labeled data in the dataset, i.e. 0.8% and 5%. Those ratios correspond to 4 and 25 labeled examples per class in CIFAR100, and to 40 and 250 samples for CIFAR10.
Table 1: Comparison of our JDCL with non-generative distillation-based techniques and generative rehearsal methods, including feature replay techniques that we mark in gray color. Results of competing methods from [9, 65].
Method CIFAR-10 CIFAR-100 ImageNet100
$T=5$ $T=5$ $T=10$ $T=5$
Joint 93.14 $\pm$ 0.16 72.32 $\pm$ 0.24 66.85 $\pm$ 2.25
Continual Joint 86.41 $\pm$ 0.32 73.07 $\pm$ 0.01 64.15 $\pm$ 0.98 50.59 $\pm$ 0.35
Fine-tuning 18.95 $\pm$ 0.20 16.92 $\pm$ 0.03 9.12 $\pm$ 0.04 13.49 $\pm$ 0.18
EWC 20.08 $\pm$ 0.51 22.28 $\pm$ 1.03 13.95 $\pm$ 1.01 21.35 $\pm$ 0.26
GKD (LWF) 49.60 $\pm$ 0.42 37.46 $\pm$ 1.18 27.91 $\pm$ 0.47 33.90 $\pm$ 0.25
MKD (iCaRL) 50.37 $\pm$ 2.15 34.01 $\pm$ 0.69 25.95 $\pm$ 0.50 33.03 $\pm$ 0.23
TKD (SS-IL) 50.46 $\pm$ 1.96 38.47 $\pm$ 0.83 28.96 $\pm$ 0.19 35.22 $\pm$ 0.22
ANCL 44.70 $\pm$ 3.24 29.46 $\pm$ 0.28 18.76 $\pm$ 0.95 29.40 $\pm$ 0.23
DGR VAE 28.23 $\pm$ 3.84 19.66 $\pm$ 0.27 10.04 $\pm$ 0.17 9.54 $\pm$ 0.26
DGR+distill 27.83 $\pm$ 1.20 21.38 $\pm$ 0.61 13.94 $\pm$ 0.13 11.77 $\pm$ 0.47
RTF 30.36 $\pm$ 1.40 17.45 $\pm$ 0.28 12.80 $\pm$ 0.78 8.03 $\pm$ 0.05
MeRGAN 51.65 $\pm$ 0.40 09.65 $\pm$ 0.14 12.34 $\pm$ 0.15 -
BIR 36.41 $\pm$ 0.82 21.75 $\pm$ 0.08 15.26 $\pm$ 0.49 8.63 $\pm$ 0.19
GFR 26.70 $\pm$ 1.90 34.80 $\pm$ 0.26 21.90 $\pm$ 0.14 32.95 $\pm$ 0.35
DDGR 43.69 $\pm$ 2.60 28.11 $\pm$ 2.58 15.99 $\pm$ 1.08 25.59 $\pm$ 2.29
DGR diffusion 59.00 $\pm$ 0.57 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 23.92 $\pm$ 0.92
GUIDE 64.47 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 39.07 $\pm$ 1.37
JDCL 83.69 $\pm$ 1.44 47.95 $\pm$ 0.61 29.04 $\pm$ 0.41 54.53 $\pm$ 2.15
Implementation details
We use the same joint diffusion model architecture in all experiments on CIFAR10 and CIFAR100, and an upscaled version of it for ImageNet100 experiments to adjust to the bigger dataset resolution. For the UNet, we follow the implementation provided in [14]. For the classifier, we use a 2-layer feed-forward net with a hidden layer of 1024 and a LeakyRELU activation function. The detailed description of the training hyperparameters is provided in the Appendix, and in the code repository [^2] https://github.com/pskiers/Joint-Diffusion-in-Latent-Space.
5.2 Main results
To compare our method with other generative rehearsal approaches, we measure its performance on several CL benchmarks. In Tab. 1, we show that our method outperforms other evaluated generative replay techniques in terms of average accuracy after the last task. On CIFAR10 and ImageNet100, our method approaches the soft upper limit, calculated as a continual joint method trained with an infinitely long replay buffer. On those benchmarks, we surpass the previous state-of-the-art score by over 19 and 15 points, which is an increase of around 30% and 40% for CIFAR10 and ImageNet100 respectively.
Figure 3: Accuracy on each task after each phase of incremental training on CIFAR100 with 5 tasks.
To better understand the advantages of JDCL over the competing methods, we perform a detailed comparison of our method against the GFR and GUIDE. To that end, in Fig. 3 we present the accuracy of each approach, on each task after each training phase on CIFAR100 with 5 tasks. JDCL significantly outperforms GUIDE in terms of knowledge retention from preceding tasks. This is especially visible when we examine the change in performance of the methods on the older tasks. Whereas GUIDE’s knowledge obtained during the first task is quickly forgotten and is almost completely lost at the end of the training, JDCL is able to maintain a good accuracy on the first task for the whole training process. Notably, the performance degradation in our method seems to be related rather to the increasing complexity of the problem instead of forgetting, as we do not see fast degradation on the oldest tasks. On the other hand, when comparing to the GFR, we can observe that our method, though worse at preserving knowledge, is much better at assimilating new information. This phenomenon can be attributed to the significantly limited plasticity of the GFR that slows down the training of its feature extractor and restricts its update to the feature distillation loss. Consequently, the model has a substantially worse final performance.
Table 2: Average accuracy with a standard deviation of different methods tested with 5-task CIFAR-10 and 10-task CIFAR-100 semi-supervised settings. The number in brackets indicates the size of the memory buffer for the labeled data. Results of competing approaches from [26]
Method CIFAR-10 CIFAR-100
0.8% labeled 5% labeled 0.8% labeled 5% labeled
[…]
Semi-supervised continual learning
Apart from the standard supervised setup, we evaluate our method on a set of semi-supervised continual scenarios and compare it to the existing approaches. As shown in Tab. 2 JDCL outperforms other methods on all CIFAR10 setups, while also achieving comparable accuracy on CIFAR100. It is worth noting that JDCL is able to reach state-of-the-art performance on CIFAR10 without a memory buffer, while also in a majority of tasks sampling fewer examples from the diffusion model than the buffer size in the related works. For CIFAR100, we believe that the reduced performance of our method is caused by an extremely weak signal from the classifier, which introduces an imbalance in our joint training. To further explain the issue, we visualize how JDCL builds latent representations in the SSL setup in the appendix.
5.3 Additional Experiments
Joint modeling Knowledge distillation 2-stage training Accuracy
83.7
68.4
48.2
36.7
32.7
The interplay between shared parametrization, distillation, and two-stage training
In this work, we propose a new generative replay method for continual learning. Apart from the main idea of our approach, which is joint diffusion and classification modeling, we combine it with two important techniques: knowledge distillation and two-stage training. In this section, we ablate the impact of each component. Notably, for experiments without joint modeling, we use a separate diffusion model with the same architecture as the joint diffusion model, and ResNet-18 as a classifier. We train and distill them in the same manner as the joint model. In Tab. 3 we show the results for the model trained without individual components. As visible, high performance of JDCL is attributed to the synergy of the proposed joint modeling, knowledge distillation technique, and two-stage training. To further highlight the interplay between those three components, we perform additional experiments.
First, we emphasize the importance of joint diffusion for effective knowledge distillation. To that end, we train a joint diffusion model and a separate classifier and diffusion on a single task. In Fig. 4, we compare the classifier’s logits distributions for a single class from the real data and rehearsal samples generated by the diffusion model trained either separately or jointly with the classifier. As visible, generations from the joint model closely match the real data logits distribution. However, for the separate classifier, the distributions diverge significantly. This discrepancy means that when updating the global model via knowledge distillation, the rehearsal data remains in-distribution for the joint classifier while being OOD for the separate classifier, which explains the observed performance gap.
Figure 4: Difference between the distribution of logit value corresponding to class 1 from CIFAR-10 for real and synthetic data. Samples generated by the joint model closely match the real data’s logits distribution, whereas those from the separately trained classifier diverge significantly.
Figure 5: Mean accuracy after each task for the ablation methods (left) and accuracy of JDCL on current and the first task with and without two-stage training (right)
To further highlight this synergy, in Fig 5 (left) we present how the average accuracy changes after each task for the ablated models. We observe that knowledge distillation and two-stage training can largely mitigate the absence of joint modeling on a task-to-task basis. However, in the CL setup, the performance differences steadily accumulate, leading to significantly worse final results.
Finally, we show that the use of two-stage training with distillation significantly stabilizes the alignment of new and old knowledge. In Fig. 5 (right) we plot the accuracy on the newly observed task, and the remaining accuracy on the first one. We can observe that although two-stage training cannot achieve as high plasticity, stable distillation from a local copy of a model to the global one leads to significantly reduced forgetting in the first task.
JDCL as a continual representation learner
In this section, we show that our method is able to learn meaningful data representation in a continual setup. We follow the evaluation approach common in self-supervised continual learning techniques, as presented in [16]. In particular, we report the accuracy of a logistic regressor trained on the features provided by our model i.e. on the set of representations of the UNet model, pooled into a single vector. In Fig. 6 we compare our model with supervised and self-supervised continual methods. We can observe that the representations learned by our model are indeed very informative. Moreover, we find that JDCL outperforms other approaches on the CIFAR100 setup with 5 tasks.
Figure 6: Accuracy of a logistic regressor trained on representations learned with different supervised and self-supervised methods on CIFAR100. Results of competing approaches from [16, 41]
Method CIFAR-10/5 CIFAR-100/5
FID Precision Recall FID Precision Recall
DGR 49.56 0.54 0.34 50.36 0.48 0.30
JDCL 32.12 0.60 0.43 39.93 0.71 0.27
Incremental generative setup comparison
Generative-replay-based continual learning methods rely heavily on the underlying generative model to preserve knowledge about the previous tasks. The quality of the samples generated by the method is therefore of utmost importance. In this section, we perform a comparison of JDCL and the diffusion model trained continually on CIFAR10 and CIFAR100, both split into 5 tasks. We present the achieved FID, Precision, and Recall after the last task in Table 4. JDCL outperforms the baseline method on all but one metric. We can observe significant improvement in the precision of generations, which results in generated rehearsal samples being clearly assigned to one of the previous classes, bringing less noisy signals to the model. As argued in [19], this trait can be the reason for the improved performance of the continually trained classifier.
Runtime analysis of JDCL for CIFAR-10/5
Method Time [GPU-hours] Accuracy
DDGR $41.29$ 43.69
DGR diffusion 28.02 59.00
GUIDE 30.60 64.47
JDCL (fast) 27.21 78.10
JDCL 53.83 83.69
6 Conclusions
Acknowledgments
This work is supported by National Centre of Science (NCP, Poland) Grants No. 2022/45/B/ST6/02817, and 2023/51/B/ST6/03004. We acknowledge PLGrid for providing computer facilities under grants no. PLG/2025/018424 and PLG/2025/018551.
Supplementary Material
Appendix A Details on initial experiment - impact of the generative replay
In this section, we describe the experimental details of the initial experiment presented in Fig 1. For our classifier, we use the ResNet152 architecture with 58.3M parameters, which is comparable to 56.3M parameters of the joint diffusion model. During the experiment, we first train both the classifier and joint diffusion model on the CIFAR10 dataset and save both models. We then use the trained joint diffusion to sample a synthetic dataset - notably, both images and labels are generated by the joint diffusion model. Then, in the first variant of the experiment, we continue training both saved models on this synthetic dataset. In a second variant, we additionally incorporate a knowledge distillation loss during the continued training. For the joint diffusion model, this involves applying the distillation approach detailed in our method, leveraging the saved model trained on the original CIFAR10 data. As for the ResNet152, we utilize the saved classifier trained on the original data and apply only the classification component of the distillation loss of our method.
Appendix B Main experiments - implementation details
We train diffusion using linear noise schedulers over 1000 steps and use DDIM with 250 steps to generate replay samples. In particular, we generate 1000, 400, and 1040 samples per class in CIFAR10, CIFAR100, and ImageNet100 respectively. For the initial local model, we train for 50,000 steps on CIFAR10, 100,000 steps on CIFAR100, and 120,000 steps on ImageNet100. Subsequent tasks are trained for 30,000 steps on CIFAR10, 70,000 steps on CIFAR100, and 90,000 steps on ImageNet100. Additionally, following [12], for the first 10000 steps of every global model training, we do not calculate the classification loss and optimize only the generative objective. We use a single AdamW optimizer with a learning rate set to 0.0002 and a batch size of 256. For all experiments, we set the classification loss scales $\alpha$ and $\alpha_{KD}$ to 0.001, and the $\beta$ hyperparameter that scales joint-diffusion loss to $0.01$.
We define three sets of data augmentations. The first one consists of flipping, resized cropping, and normalization to range [1,1][-1,1], and is used to train the denoising model. The second, strong augmentation set includes flipping, cropping, the RandAugment [8] augmentation policy, and normalization to range [1,1][-1,1]. The third and final, weak augmentation set consists of flipping, cropping, and normalization to range [1,1][-1,1]. We use the strong augmentation set to perturb the images for the classifier in the supervised setups.
Additionally, following [63], we use our strong and weak augmentation sets to compute the semi-supervised loss. We also set the $\tau$ threshold to $0.95$ and we set the supervised-to-unsupervised batch size ratio to 7, which corresponds to a supervised batch size of 32 and an unsupervised batch size of 224.
Appendix C Stability of the JDCL’s representation space in an incremental setup
Figure 7: Accuracy of the kNN-clasifier trained on representations provided by the model after the first task of class-incremental CIFAR100 setup with 5 tasks
One of the most important features of our methods is its ability to learn a meaningful representation space shared by the generative and discriminative parts of the model. To better understand how this representation space changes during the class-incremental training, we propose to measure the drift of representations, with the following experiment on a 5-task CIFAR100 setup. We first extract data representations from two parts of our UNet model, the H-space inside of the generative part of our model, and the penultimate layer of the classifier head. We train a kNN-classifier on top of those representations provided by the model after the first task. Then, we evaluate how the accuracy of this classifier changes when the representations for data from the first task are extracted with models trained incrementally on later tasks. We present the results in Fig. 7. We observe that the representations extracted from the UNet change significantly slower than those from the classifier. This means that representations that are not used for classification drift significantly slower than those that are classifier-specific, which highlights the benefits of joint modeling in a continual setup.
Figure 8: Visualization of the representation spaces (umap) of different methods trained on a partially labeled dataset. Labeled data samples are marked with an "X", while unlabeled data points are marked with circles. All the models were trained on the first 10 classes of CIFAR100 with 5% of the data being labeled.